Find forces in members DF, DE, and EC and state whether they are in tension or compression

12 ft 12 ft 12 ft 12 ft 12 ft 12 ft H 9 ft 9 ft D G K 9 ft В. E 1 M 9 ft) C L А 4 kips 4 kips QN
Answer #1

Keep in mind that negative sign indicates opposite direction of forces in member from assumed direction.Jolution: FBD : This type of problem, method of section is easy. 12 ft I + F I G D E k J 5 -B aft M L A N Ax 4 kips 4 kip AySD . To considor left portion, dxk ng must be known To find dy & mo EM20 = Rev X (1286)=0 IRN:o 199 Fyr Ax+4-4 = 0 Ax = 0 EFy- 10 kips (compressim FCE = 10 kips (compression) EME=0 – FDF X 7.2 + 4% E P =D (- EP= 9+9=18) FDF87.2 = 18x4 For = to kip (T